OpenStudy (anonymous):

Multiply (x^6+x^5)(x^6+x^5-x^4)

OpenStudy (unklerhaukus):

\[ (x^6+x^5)(x^6+x^5-x^4)=x^5(x+1)x^4(x^2+x-1) \] \[=x^9(x+1)(x^2+x-1)\]